OpenStudy (firejay5):

Find the value of each variable. If your answer is NOT an integer, leave it in simplest radical form. Show work!!! Medal will be rewarded for patronage! :D

OpenStudy (firejay5):



OpenStudy (firejay5):

@jdoe0001 @whpalmer4 @jim_thompson5910

OpenStudy (whpalmer4):

are the two upper sides of equal length?

OpenStudy (firejay5):

@whpalmer4 I am confused on your question

OpenStudy (whpalmer4):

the triangle has 3 sides. the bottom one is labeled \(9\sqrt{2}\). The right one is labeled \(a\). Is the left one of a known length?

OpenStudy (firejay5):

no

OpenStudy (whpalmer4):

Are you sure? You can't figure it out from what you know about that triangle? If you can't, I'm not sure you're going to have much luck with this problem unless you know trigonometry.

OpenStudy (whpalmer4):

What kind of a triangle is that?

OpenStudy (whpalmer4):

@firejay5 My 10 year old correctly identified that as an isosceles triangle and told me the length of the upper left side. Maybe you should have another look, eh?

OpenStudy (whpalmer4):

and now he's figured out \(a\). I had to remind him of the formula for the Pythagorean theorem, so I'll do the same for you:\[a^2+b^2=c^2\]\(a,b\) are the lengths of the sides, \(c\) is the length of the hypotenuse. Don't tag me until you've gotten the correct answer here.

OpenStudy (firejay5):

I got A = 9

OpenStudy (whpalmer4):

Great! Your tagging privileges are reinstated :-)

OpenStudy (firejay5):

what do you mean by that?